OpenStudy (anonymous):

use the half angle formula to find the exact value of 165

myininaya (myininaya):

165 is equal to 165 No half angle formula needed.

myininaya (myininaya):

I think your question is missing something honestly...

OpenStudy (anonymous):

how do I find the exact values of sin, cos and tang?

myininaya (myininaya):

sin(165)=\[\sin(\frac{330}{2})=\sqrt{\frac{1}{2}(1-\cos(330))}\] Now you simplify.

myininaya (myininaya):

\[\sin(\frac{x}{2})=\pm \sqrt{\frac{1}{2}(1-\cos(x))} \text{ and } \cos(\frac{x}{2})=\pm \sqrt{\frac{1}{2}(1+\cos(x))}\] since the angle is 165 we are in between 90 and 180 which means sin is positive and cos is negative

OpenStudy (anonymous):

Thank you so much